OpenStudy (anonymous):

find the distance between points (-2,2) and (-5,4)

OpenStudy (experimentx):

sqrt(9+4) = sqrt(13)

OpenStudy (anonymous):

sqrt(9+4) = sqrt(13) probably

OpenStudy (anonymous):

Oh, Rohan you're 13 too ...just 13

OpenStudy (anonymous):

yaa

OpenStudy (callisto):

Use distance formula \[\sqrt{(x _{2}-x_{1})^2 +(y _{2}-y _{1})^2}\]\[=\sqrt{[(-2)-(-5)])^2 +(2-4)^2}\]\[=\sqrt{3^2 +(-2)^2}\]\[=\sqrt{9 +4}\][=\sqrt{13}\]

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

\[=\sqrt{13}\]

OpenStudy (callisto):

last line is \[=\sqrt{13}\]

OpenStudy (callisto):

Thanks @satellite73